### v1

# High-Fidelity Data-Driven Dynamics Model for Reinforcement Learning-based Magnetic Control in HL-3 Tokamak

Abstract
The drive to control tokamaks, a prominent technology in nuclear fusion, is essential due to its potential to provide a virtually unlimited source of clean energy. Reinforcement learning (RL) promises improved flexibility to manage the intricate and non-linear dynamics of the plasma encapsulated in a tokamak. However, RL typically requires substantial interaction with a simulator capable of accurately evolving the high-dimensional plasma state. Compared to first-principle-based simulators, whose intense computations lead to sluggish RL training, we devise an effective method to acquire a fully data-driven simulator, by mitigating the arising compounding error issue due to the underlying autoregressive nature. With high accuracy and appealing extrapolation capability, this high-fidelity dynamics model subsequently enables the rapid training of a qualified RL agent to directly generate engineering-reasonable magnetic coil commands, aiming at the desired long-term targets of plasma current and last closed flux surface. Together with a surrogate magnetic equilibrium reconstruction model EFITNN, the RL agent successfully maintains a $100$-ms, $1$ kHz trajectory control with accurate waveform tracking on the HL-3 tokamak. Furthermore, it also demonstrates the feasibility of zero-shot adaptation to changed triangularity targets, confirming the robustness of the developed data-driven dynamics model. Our work underscores the advantage of fully data-driven dynamics models in yielding RL-based trajectory control policies at a sufficiently fast pace, an anticipated engineering requirement in daily discharge practices for the upcoming ITER device.
Keywords: data-driven, dynamics model, reinforcement learning, magnetic control, HL-3 tokamak
1 Introduction
The objective of the tokamak, a torus-shaped nuclear fusion device, is to provide sustainable energy by effectively confining high-temperature plasmas using magnetic fields. The stable confinement of plasma in tokamak depends significantly on well-calibrated control approaches, which can be developed through a thorough understanding of the underlying dynamics [1, 2, 3]. Rather than providing an exhaustive interpretation of fundamental physics, control-oriented dynamics models deliver significant advantages in simplicity with reasonable accuracy [4], assisting the precise manipulation of various actuators toward desired configurations. For instance, the reliable prediction of rapidly growing plasma instabilities contributes to proactively suppressing disruptions [5, 6, 7, 8], while an accurate long-term evolution of the plasma dynamics supports Reinforcement Learning (RL)-based controller to meticulously regulate high-dimensional, mutually coupled magnetic coils towards target plasma current and shape [9, 10].
Classically, a high-fidelity, autoregression-capable dynamics model [11, 12] can be built on first-principles-based computations. A notable example is the Forward Grad–Shafranov Evolutive (FGE) model, which solves the free-boundary equilibrium while accounting for the plasma current evolution coupled to controller dynamics [13]. Through accurately synthesizing high-dimensional magnetic flux and local magnetic field measurements as outputs of the tokamak sensors, this model enables the offline training of RL-based magnetic controller [9]. Although this physics-driven approach can give highly detailed simulations, an over-reliance on such simulators without incorporating sufficient real-world data may reduce the generalizability of RL policies [14]. Additionally, each training step has to await the FGE simulator for the cumbersome computations of lumped-parameter equations [9], making the training of RL policies possibly take hours [10]. In contrast, ITER, the International Thermonuclear Experimental Reactor, is expected to be capable of executing the next pulse within a minimum of $30$ minutes [15]. Despite progress to accelerate training, the physics-driven simulator is still not competent for training RL-based control policies at a sufficiently fast pace [10], an anticipated engineering requirement in daily discharge practices for ITER.
Compared to first-principles-based approaches, historical discharge data offers the potential to develop models with significantly faster speed. For example, experimental data facilitates the adoption of linearized system identification techniques to approximately reproduce the response of coupled plasma profiles under variations of given actuators [16, 14]. Nevertheless, the validity of the linearized dynamics model is limited to the vicinity of the plasma equilibrium that needs to be tracked [17]. An alternative approach resorts to Artificial Intelligence (AI) models, as Deep Neural Network (DNN)-based models demonstrate satisfactory accuracy [18, 19, 20, 21, 22] with preliminary dynamics modeling results [12, 11, 5]. Prominently, a DNN model that competently forecasts the future tearing likelihood from various plasma profiles and tokamak actuations, has successfully contributed to the prediction-driven planning of a high-level policy to yield desired beam power and plasma triangularity target for tearing avoidance [5]. Though such a process disentangles the predictive power of the dynamics model from the policy learning procedure, building a high-fidelity, data-driven simulator for training a high-dimensional, high-frequency RL controller remains challenging. Unless optimizations are adopted to make long-range predictions, the autoregressive nature of such simulators can lead to compounding errors, which gradually accumulate over long-term multiple-step evolution, resulting in significant discrepancies and potentially catastrophic outcomes [23].
[…]
Fig. 1: The overall architecture of the plasma current and shape control in the HL-3 tokamak. a, The deployment of PID and RL control systems on HL-3. The historical interactions between the PID and the HL-3 tokamak produce the dataset for learning the data-driven dynamics model, which serves as an environment for RL training. RL controller is tasked with controlling the plasma current ($I_{p}$) and six shape parameters (minor radius $a$, elongation $\kappa$, upper triangularity $\delta_{u}$, lower triangularity $\delta_{l}$, radial position $R$ and vertical position $Z$ of plasma geometric center), with the output actions consisting of voltage commands for $17$ magnetic coils of the HL-3 tokamak. b, Schematic diagram of the localized model predict planning, wherein the dynamics model refrains from autoregressive extension and guides responsive actions only. c, Schematic diagram of long-term trajectory control. Different from using the physics-based simulator in [9], our work utilizes the dynamics model autoregressively to facilitate training RL-based controller. d, The proposed training diagram of the dynamics model, including scheduling sampling, multi-task adaptive weight adjustment, self-attention, and model ensembling.
To combat the long-term divergence issue, one intuitive means is to unroll the single-step prediction to incorporate multi-prediction losses during training [24, 25, 26, 27, 28]; while another line of work lies in implicitly [29, 30, 31] or explicitly [32, 33] learning a specific dynamics model for every $n$-step prediction. Albeit the practicability in low-dimensional scenarios [25, 26, 27, 28, 29, 30, 31, 32], both strategies imply significant training difficulties with limited scalability for a magnetic control that usually lasts for hundreds or even thousands of steps. Therefore, on top of single-step prediction, several techniques like ensembling [34, 35], principle component analysis [36] and noise layer [12, 11] have been leveraged to ameliorate this issue. However, to our best knowledge, the data-driven dynamics model is still far from perfect for long-term, high-dimensional trajectory control. Inevitably, Proportional-Integral-Derivative (PID) controllers [12, 11, 5], historically logged multi-variate commands [36, 37], or online Model Predictive Control (MPC) [35] indispensably complements for achieving localized short-term, low-level magnetic control. Nevertheless, PID controllers imply tremendous engineering and design efforts before deployment [9], while online MPC suffers from the cumbersome online autoregression of the dynamics model for trajectory sampling [6, 35]. Consequently, due to the rare readiness of either low-level controllers or physical simulators, the limited accuracy with only short-term guarantees in existing data-driven dynamic models hinders the learning of high-dimensional, long-term trajectory control policies.
In this paper, we aim to provide a data-driven dynamics model, that exclusively learns from the discharge logs collected through the interaction between PID controllers and the tokamak device. The dynamics model is designed to efficiently support the fast training and practical deployment of a long-term, low-level RL-based agent, capable of directly generating voltage commands for up to $17$ magnetic coils. The system architecture is shown in Fig. 1. As shown in Fig. 1a, EFITNN [38], a surrogate model of magnetic equilibrium reconstruction code EFIT [39, 40], is leveraged to meaningfully transform the high-dimensional magnetic measurements to lower-dimensional plasma shape information while maintaining the physical interpretability and supporting real-time deployment. Subsequently, a Long Short-Term Memory Network (LSTM)-based [41] dynamics model is developed to track the evolution of plasma shape and current, based on accessible zero-dimensional plasma parameters and executed actuations. Benefiting from effectively leveraging several useful techniques including scheduling sampling [42, 43], adaptive weight adjustment [44, 45], self-attention [46] and model ensembling [47], the trained dynamics model demonstrates exceptional generalization capability on unseen plasma current regions during training and sensitively captures shape variations in typical scenarios, such as transitioning from limiter to (advanced) divertor configuration [48], and soft ramp-down [49] during discharges. Unlike localized short-term planning [5, 12, 11] shown in Fig. 1b, this data-driven dynamics model enables fully RL-based $100$-ms, $1$-kHz low-level trajectory control, achieving the desired Last Closed Flux Surface (LCFS) [50] via six shape and position parameters; while as in Fig. 1c, the training of $250,000$ iterations completes within $22$ minutes on the customer-grade Nvidia GeForce RTX 4090 GPU, fully satisfying the anticipation for RL-based magnetic control in ITER. It even achieves zero-shot triangularity adjustment, validating the effectiveness and reliability of the data-driven dynamics model.
2 Results
2.1 Construction of Data-Driven Dynamics Model
Fig. 2: The demonstration case to predict plasma shape and position parameters for shot #6065 with and without the involvement of Scheduled Sampling (SS). Compared with actual discharge data (black solid line), a model trained with scheduled sampling (red dashed line) more effectively mitigates the compounding error than that without scheduled sampling (blue dashed line). The shaded area represents the results of ensembling $5$ independent models.
To train a $1$-kHz RL-based control policy, it is essential to build a high-fidelity data-driven dynamics model, which can autoregressively produce accurate predictions for up to $25$ variates. These predictive variates encompass the plasma current, $6$ plasma shape and position parameters, and $18$ channels of feedforward coil currents, which together correspond to a $44$-dimensional indispensable input (Table 2) and support the training of an engineering-reasonable policy for controlling $17$ magnetic coil voltages. In comparison to yielding $100$ ms-averaged evolution [12], forecasting single-variate tearability [5], or only tuning the total power and torque injected from the neutral beams [36], the significantly higher resolution and a larger number of coupling actuators implies to pose considerable challenges and requires a more comprehensive design.
During training, to mitigate the compounding errors, scheduled sampling, which blends the historical logged data and model output [42, 43], is leveraged. Fig. 2 shows this easy-to-implement technique ensures the training scalability and brings appealing results; in Methods, a more thorough comparison with alternative solutions such as noise layer [12] shows the flexibility of this training approach. On this top, training the dynamics model incorporates several useful techniques. In particular, adaptive weight adjustment [44, 45], which simplifies the tuning of weights for combining multi-task loss functions [45] and guarantees resilient and steady model learning [44], is adopted to capture the implicit relationship among plasma shape and position parameters. Meanwhile, auxiliary techniques, such as self-attention [46] and model ensembling [47], further enhance accuracy through a concerted effort. Fig. 3a presents the ablation study in terms of Mean Absolute Error (MAE) and demonstrates the incremental contribution of each technique to train the dynamics model. Besides, the model maintains high consistency across different tasks, achieving at least a $58.8\%$ reduction in MAE than the standard LSTM-based training. In addition, Fig. 3b provides the statistical measure in terms of the coefficient of determination $R^{2}$ and shows the autoregressive output of the dynamics model can fit the actual data well.
[…]
Fig. 3: Accuracy of the dynamics model. a, Ablations for key techniques in terms of MAE. SS: Scheduled Sampling, EN: Ensemble, SA: Self-Attention, AW: Adaptive Weight. Each group depicted here is derived by incorporating the corresponding module onto the basis of the preceding group. For instance, the notation “+ EN” signifies the integration of the Ensemble into the LSTM model that already incorporates scheduled sampling (“+ SS”). b, Goodness-of-fit of the dynamics model in terms of the coefficient of determination $R^{2}$.
2.2 Prediction Accuracy under Abrupt Shape variations
[…]
Fig. 4: The accuracy of dynamics model to predict abrupt plasma shape variations. a, Shot #6221, an H-mode discharge with ELMs and a plasma current of $300$ kA. b, Shot #6063, an advanced divertor configuration discharge with a plasma current of $500$ kA, with advanced divertor configuration undergoing a complete transformation between $1300$ ms and $1500$ ms. c, Shot #6232, an ITB discharge with a plasma current of $600$ kA. d, The LCFS variations correspond to the transition from limiter to divertor shape and soft ramp-down phases for shot #6221 ($300$ kA) and shot #6232 ($600$ kA), respectively.
In the previous section, we demonstrated the capability of the data-driven dynamics model in learning the high-dimensional evolution. In this section, we switch to investigate the prediction accuracy under abrupt plasma shape variations, which are commonly encountered due to the transition from limiter to divertor shape, the advanced divertor configuration transformation, or the initialization of soft ramp-down. Taking the example of shot #6221, an H-mode discharge with edge localized modes (ELMs), Fig. 4d manifests the significant impact of these activities on the plasma LCFS. Specifically, following the shift from the limiter to the divertor configuration, the plasma’s elongation $\kappa$, upper triangularity $\delta_{u}$, and lower triangularity $\delta_{l}$ surge abruptly. Fortunately, as shown in Fig. 4a, the dynamics model sensitively captures this variation. Similar robustness in predicting significant fluctuations in elongation and horizontal displacement can be observed in the initial segment of the ramp-down phase, where magnetic control still dominates despite the plasma’s strong interaction with the wall producing significant non-magnetic effects. On the other hand, Fig. 4b and Fig. 4c validate the accuracy of the dynamics model to predict abrupt shape variations in both advanced divertor configuration and Internal Transport Barrier (ITB) discharges [51, 52].
2.3 Extrapolation Capability on Unseen Plasma Current
Fig. 5: MAE averaging over 30 shots of 700 kA discharges. The numbers in the brackets correspond to the number of $700$ kA discharges included in the training dataset.
Fig. 6: Test results of shot #4042 under the dynamics model from zero-shot (blue dashed line), $2$-shot (green dashdot line), and full-shot training (red dotted line). The shaded area represents the results of the model ensemble, while actual discharge data is given in a black solid line.
Consistent with the staged research plan in ITER to gradually increase plasma current ($I_{p}$) [15], it is crucial to assess the extrapolation capability of a dynamics model trained on lower-$I_{p}$ datasets when applied to higher $I_{p}$ plasma discharge scenario. Therefore, we investigate the marginal contributions of incorporating various numbers of additional $700$ kA discharges to fine-tune a dynamics model initially trained on data with plasma currents ranging from $300$ to $600$ kA.
Fig. 5 presents the corresponding results averaging over a testing dataset consisting of $30$ shots of $700$ kA discharges, where the zero-shot training implies no involvement of data from $700$ kA discharges, while full-shot learning indicates using data from all available $67$ shots. As depicted in Fig. 6, under zero-shot training, the model gives a $152.6$ kA discrepancy and fails to reach the maximum plasma current value of $700$ kA, highlighting the generalization difficulty to higher currents without any prior information. Nevertheless, with the introduction of just $2$ or $4$ additional shots of $700$ kA data for training, the prediction accuracy improves significantly, with a $90\%$ reduction in MAE. Meanwhile, both Fig. 5 and Fig. 6 indicate that with the inclusion of $6$ additional shots, the performance closely approaches that of full-shot training, demonstrating the model’s strong extrapolation capability even with a limited dataset. On the other hand, for the geometric parameters in the $700$ kA discharge, a zero-shot model maintains an acceptably low prediction error, suggesting more robust extrapolative abilities for shape and position parameters. These experiments carry encouraging implications for both the predictive and control aspects of staged discharge within the upcoming ITER device.
2.4 RL-based On-Device Control Results
[…]
Fig. 7: Target tracking control results of RL on both simulation (a) and actual devices (b). a, Simulating control results of plasma current $I_{p}$, shape and position parameters $a$, $\kappa$, $\delta_{u}$, $\delta_{l}$, $R$, and $Z$ are achieved, with the handover time set at the $1500$ ms. b, in shot #6691, the experimental data (red solid line) and the post-experiment replay of the dynamics model (blue dashdot line), as well as the target (black dashed line).
In line with the long-term trajectory magnetic control in [9], we aim to learn an RL-based policy to generate direct coil voltage commands for target plasma shape. This policy is trained through interactions with an autoregressively evolving dynamics model. During training, an actor-critic-based Proximal Policy Optimization (PPO) [53] is leveraged for learning to achieve the desired plasma current and shape, which is identical to the discharge result between $1500$ and $2000$ ms of shot #6226 of the HL-3 tokamak, an H-mode discharge with ELMs and a plasma current of $300$ kA. In PPO [53], the “actor” utilizes a policy to generate physics-reasonable magnetic coil voltage commands (see Methods) and modulate the plasma to the configuration of interest, while the “critic” evaluates the effectiveness of the current policy by estimating state values or advantages for potential policy refinement. Fig. 7a demonstrates the target tracking accuracy of the well-trained PPO actor within $1500$ $2000$ ms’ simulations. The control error of the plasma current is $2.55$ kA, or equivalently $0.85$% of the target for a $300$ kA discharge. The minor radius control error is $0.34$ cm, which corresponds to $0.52$% of the $65$ cm minor radius characteristic of the HL-3 tokamak device. Additionally, the control errors for the geometric center coordinates, both $R$ and $Z$, are constrained within a maximum of $1$ cm. The results indicate that the RL agent, supported by a data-driven dynamics model, can effectively explore and assimilate strategies that are advantageous for achieving discharge control objectives.
Inspired by these promising results, we further evaluate the performance of the RL-based policy by integrating the well-trained PPO actor with the Plasma Control System (PCS) of the HL-3 tokamak. Consistent with [9], experiments are executed without further tuning weights of the PPO actor after training. In other words, there is a “zero-shot” transfer from the simulation to the real device. Meanwhile, to provide a real-time plasma shape diagnostic and complete the shape-control feedback loop, we utilize our previously developed EFITNN [38], a surrogate model of the EFIT code [39, 40] consisting of several fully connected neural network layers with residual connections [54]. In comparison with offline EFIT, EFITNN yields a relative error of less than $1\%$ and consumes just $0.08$ ms on NVIDIA A100 Tensor Core GPU with TensorRT. The adoption of data-driven EFITNN contributes to effectively reducing the learning expenses of the dynamics model and facilitating real-time control. A summarized comparison between EFITNN and offline EFIT is presented in Fig. 8.
Fig. 8: Error between EFITNN and offline EFIT.
2.4.1 RL-based $100$-ms Magnetic Control for Target Tracking
To thoroughly evaluate the flexibility of RL-based target tracking, we adopt a progressive approach. Initially, we assess the agent’s target tracking ability and ensure it aligns with the simulations. For this purpose, the experimental target remains identical to the reference shot #6226 for PPO training and the PPO actor responds to the variations in plasma shape and position parameters diagnosed by EFITNN. Despite the consistency in targets, noticeable differences arise during the PID control due to the inherent uncertainties in tokamak discharge and simultaneously coupled operations for HL-3 maintenance and testing.
In shot #6691, alongside the traditional PID control, we conduct an RL-based magnetic control from $1600$ ms to $1700$ ms. During this period, a $10$-ms transition period is established to ensure a stable switch between both control modes. Fig. 7b provides the corresponding control results.
Throughout the $100$-ms RL control, the minor radius $a$, elongation $\kappa$, and lower triangularity $\delta_{l}$ all progress with fast correction and accurate results as expected. However, due to the excessive horizontal displacement and acceleration on the handover time, radial position $R$ is not corrected fully, despite the adjustments being in the proper direction. Similar observations apply to the upper triangularity $\delta_{u}$, which deviates rapidly before $1600$ ms and is corrected progressively during the RL-based control period. Regarding plasma current $I_{p}$ and vertical position $Z$, the measured parameters fluctuate around the targets with moderate control errors. These outcomes manifest the robustness of the on-device RL-based control. Additionally, a post-experiment replay of the dynamics model (see Methods) closely aligns with the experimental results, confirming the effective and reliable support for RL decision-making.
2.4.2 RL-based Control for Zero-Shot Adaptation to Changed Triangularity Target
[…]
Fig. 9: Zero-shot triangularity control in shot #6698. a, The upper and lower triangular control results of shot #6698. Compared to the target used for RL training (black dashed line), the target for this discharge (red dashed line) has a triangularity decreased by $0.2$. The red solid line illustrates the experimental result, and the blue dashdot line represents the post-experiment replay of our dynamics model. b, Plasma LCFS reconstructed from six plasma shape and position parameters ($a$, $\kappa$, $\delta_{l}$, $\delta_{u}$, $R$, $Z$) during the RL control phase.
Triangularity is a critical shape parameter in tokamak plasmas, significantly affecting both vertical stability and the structure of the plasma boundary [55]. Consequently, the ability to control triangularity is crucial for optimizing tokamak plasma equilibria and performance. To validate the flexibility of RL, we intentionally decrease the upper and lower triangular shape parameters of the plasma by $0.2$ from those target values during the training phase. The performance of the PPO actor on triangularity targets not encountered during training allows us to evaluate the agent’s adaptability to varying control objectives.
In shot #6698, similar to the procedure described in Section 2.4.1, RL assumes the discharge control task from PID from $1600$ ms to $1700$ ms. Fig. 9a illustrates the corresponding control results of the upper and lower triangularity. Both the lower triangularity $\delta_{l}$ and the upper triangularity $\delta_{u}$ successfully reach the new target values. Specifically, the lower triangularity $\delta_{l}$ swiftly progresses towards the target following a transition period of approximately $10$ ms, finally reaching it around $1650$ ms. Meanwhile, the upper triangularity $\delta_{u}$ follows a gradual trajectory toward the target during the initial $10$-ms transition phase, accelerates its adjustment after around $20$ ms, and aligns with the target at approximately $1650$ ms. It then maintains alignment with the target trajectory for the remaining $50$ ms. These zero-shot triangularity control results verify the versatility of RL towards effective control in previously unencountered scenarios. Furthermore, the post-experiment replay of the dynamics model from $1600$ ms to $1700$ ms reveals a consistent trend, underscoring the model’s generalizability and robustness for supporting RL in exploratory applications.
3 Discussion
We present a technique to learn a fully data-driven dynamics model to characterize the evolution of plasma in a tokamak device. The dynamics model, which leverages multiple components to mitigate the compounding error issue and improve accuracy during autoregression, exhibits high fidelity and appealing extrapolation capability. It effectively supports the fast training of an RL-based high-dimensional trajectory control policy, which together with the real-time plasma diagnostic empowered by EFITNN, can directly manipulate the magnetic coil voltages to transform the plasma state toward desired configurations.
Fig. 10: Comparison of the training and deployment between a, MPC with online, limited-term dynamics model autoregression and b, PPO interacting with dynamics model for training only.
Our work demonstrates the feasibility of data-driven, end-to-end learning and control of plasma current and shape. By avoiding reliance on first-principle-based simulators [9] and auxiliary controlling commands from historical data or controllers [12, 11, 5, 36, 37, 35], our approach significantly differs from existing works. High-dimensional control over a learned model falls into the general scope of model-based reinforcement learning (MBRL) [24]. In this regard, as in Fig. 1b, instead of evolving the plasma state, Ref. [5] mainly uses the output of the learned dynamics model to generate a competent reward function, which limits its applicability to localized reactive control only. Meanwhile, as in Fig. 10, Ref. [35] conducts a receding (limited-term) horizon control through computation-intensive MPC, which involves cumbersome online autoregression of the dynamics model for trajectory sampling, making real-time deployment challenging [56]. On the contrary, as in Fig. 1c, our work provides a proof-of-concept study on long-term high-dimensional dynamics model learning, which further enables the training and deployment of the independent PPO actor for complex magnetic control.
We also note that our work is still in the early stages and owing to limited experimental opportunities, a more comprehensive validation is yet to be carried out. Therefore, further experiments and fine-tuning can be explored for more useful applications. For example, more advanced learning and control objectives involving plasma instability and disruption avoidance [57, 58, 59] can be studied. Additionally, progress in AI fields including advanced DNN design with improved capabilities in time-varying dynamics learning [60, 61, 62, 63, 64] and decision-making [65, 66], and orthogonal ingredients from MBRL to cope with out-of-distribution (OOD) issues [67, 68, 69] can be leveraged. Besides, transfer learning across tokamak devices [70] warrants further investigation. This could verify the potential advantages of a fully data-driven method over first-principle-based simulators, which require sophisticated device adaptation. We believe that as a meaningful extension to broaden the capability of data-driven approaches in the fusion area, our approach promises to more efficiently provide qualified solutions to daily discharge operations in future tokamak devices.
4 Methods
4.1 HL-3
The HL-3 (previously named HL-2M) Tokamak, an experimental fusion device constructed at the Southwestern Institute of Physics in China, aims to advance the comprehension of fundamental physics and engineering principles crucial for the success of global fusion projects like ITER. HL-3 boasts a plasma current capacity ranging from $1.5$ to $3$ MA and operates within a magnetic field strength of $2.3$ T. Its geometric parameters consist of a major radius of $1.78$ m, a minor radius of $0.65$ m, and an aspect ratio of approximately $2.8$, reflecting its toroidal structure. As shown in Fig. 1a, the poloidal field configuration of HL-3 is intricately managed by a Central Solenoid (CS) coil, complemented by an ensemble of eight pairs of poloidal field (PF) coils strategically distributed as inner coils (PF1-PF4), top-bottom coils (PF5-PF6), and outer coils (PF7-PF8), with the latter trio playing a pivotal role in establishing the divertor configuration, hence their nomenclature as divertor coils [71, 72, 73]. In particular, HL-3 directly manipulates the voltage of $17$ coils, comprising of $1$ CS coil and $16$ PF coils. This manipulation indirectly influences the coil currents and regulates the magnetic field within the vacuum chamber, enabling the confinement and discharge of plasma.
4.2 Data Pre-Processing
Table 1: Flat-top plasma current distribution in the dataset.
Plasma Current (kA) $300$ $400$ $500$ $600$ $700$ $800$ $1000$
Number of Shot $258$ $94$ $245$ $77$ $97$ $32$ $29$
Typically, the plasma discharge process can be divided into three distinct phases (i.e., ramp-up, flat-top, and ramp-down phases). However, for the ramp-up phase’s plasma, the relatively low current and the difficulty in calculating the eddy currents in passive conductor structures (such as the vacuum chamber and the first wall) often result in inaccurate shape calculations, weakening the correlation between the coil voltage commands and plasma shape. Therefore, considering that the primary objective of the dynamics model is to capture the complex evolution of the plasma state and ultimately facilitate efficient RL training, model training is exclusively based on data from the ramp-up phase where the plasma current surpasses $100$ kA and flat-top phase with a duration of at least $500$ ms. Based on the historical discharge logs of HL-3 across campaigns from 2022 to 2024, after such a meticulous selection procedure, a dataset comprising $832$ successful discharges has been curated, with a temporal resolution of $1$ ms, totaling $1,833,516$ time slices. The categorization of these $832$ shots by flat-top plasma current is detailed in Table 1.
Table 2: Input and output variables used in dynamics model, with the length of the time series $n=30$ and the temporal resolution $\Delta{t}=1$ ms. As the CS coil is composed of two parallel connected sets of coils, the dimension of coil current ($I_{c}$) equals that of coil voltage ($U$) plus $1$.
Input variable Unit Number of channels Description Time
$I_{p}$ kA 1 Plasma current ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$a$ cm 1 Minor radius ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\kappa$ dimensionless 1 Elongation ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\delta_{u}$ dimensionless 1 Upper triangularity ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\delta_{l}$ dimensionless 1 Lower triangularity ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$R$ cm 1 Radial position of plasma geometric center ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$Z$ cm 1 Vertical position of plasma geometric center ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$I_{c}$ A 18 Coil current ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$U$ V 17 Coil voltage ($t-(n-1)\Delta{t}$$\sim$($t$)
$TF$ T 1 Toroidal magnetic field ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$Vloop$ V 1 Loop voltage ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
Output variable Unit Number of channels Description Time
$I_{p}$ kA 1 Plasma current $t$
$a$ cm 1 Minor radius $t$
$\kappa$ dimensionless 1 Elongation $t$
$\delta_{u}$ dimensionless 1 Upper triangularity $t$
$\delta_{l}$ dimensionless 1 Lower triangularity $t$
$R$ cm 1 Radial position of plasma geometric center $t$
$Z$ cm 1 Vertical position of plasma geometric center $t$
$I_{c}$ A 18 Coil current $t$
Acknowledging the importance of the historical plasma state and coil commands, the dynamics model integrates a time series as its input to better capture the plasma evolution towards equilibrium. Specifically, given the $1$-kHz control frequency, the model uses a $30$-ms historical dataset for single-step prediction of plasma current, shape and position parameters, as well as the coil currents. A summarized list of the input and output variables of the dynamics model can be found in Table 2. Briefly, this model learns up to $25$ variables related to plasma current, shape, and position, as well as coil current from a $44$-dimensional input, which includes autoregressive predictions of these $25$ variables, $17$ executed coil voltage configurations, the toroidal magnetic field, and loop voltage. In addition, $n=30$ represents the length of the time series, and $\Delta{t}=1$ ms is the temporal resolution. Notably, as the CS coil is composed of two parallel connected sets of coils, the dimension of coil current ($I_{c}$) equals that of coil voltage ($U$) plus $1$.
Throughout the discharge experiments in HL-3, engineering conditions change continuously. As a result, we utilize the most recent discharges for testing while earlier datasets are used for training. This methodology contributes to learning from updated data distributions and accommodates to latest discharge environment. To prevent model overfitting, we designate $50$ discharges from Shot #5893 to Shot #6055 as the validation set, and $70$ discharges, including Shots #6056-#6265 and #5095-#5105, as the test set. Note that the range from Shot #5095 to #5105 corresponds to $10$ discharges of advanced divertor configuration, which exhibit more pronounced morphological variations than standard discharges. During training, a $15$-ms smoothing process is implemented on the training data to effectively mitigate non-electromagnetic interference. Since the training dataset suffers from a paucity of similar data instances, the corresponding results in Fig. 3 could partially validate the accuracy and extrapolation capability of the dynamics model.
4.3 DNN Design
4.3.1 DNN Structure for Dynamics Model
[…]
Fig. 11: The testing results of response model with different structures on the same test dataset. TRF: Transformer; All: Scheduled Sampling + Ensemble + Self-Attention + Adaptive Weight.
Due to the practical demands, plasma discharges are often carried out with identical or similar parameter sets to replicate specific physical outcomes, resulting in a lack of diversity in the dataset. Therefore, although the Transformer excels in parallel computation and can apprehend holistic information owing to its self-attention mechanism [46], it faces overfitting issues under limited and narrow data distribution. The testing MAE for both the Transformer and LSTM networks, as shown in Fig. 11, suggests that under the same training methods, dynamics models utilizing LSTM networks typically outperform those based on the Transformer. As a result, we opt for LSTM as the backbone of our dynamics model.
We utilize a stack of DNNs including a three-layer LSTM network, a self-attention module, and a three-layer fully connected MultiLayer Perceptron (MLP). Specifically, each LSTM layer comprises $128$ hidden neurons with a dropout rate of $0.2$. For the $30$ consecutive steps’ $44$-dimensional input, as in Table 2, the third LSTM layer produces an intermediate hidden layer output with dimensions of $30\times 128$. Subsequently, the inner product between each $128$-length tensor and the last one is computed, followed by a softmax-based normalization operation to scale the values between 00 and $1$. These self-attention values serve as attention weights to aggregate the original LSTM outputs as a $128$-length vector. The integration of a self-attention mechanism into the LSTM allows the assignment of distinctive weights to the intermediate output, rather than placing equal emphasis and neglecting the inter-differences. Fig. 3a validates the effectiveness of this modification. Afterward, the self-attention-processed vector undergoes three MLPs with $128$, $64$, and $25$ neurons and a ReLU activation function, culminating in the single-step prediction of $25$ variables as specified in Table 2.
4.3.2 DNN Structure for EFITNN
EFITNN [38] is a surrogate model trained with input and output provided by an offline Equilibrium FIT (EFIT) code [39, 40]. It accepts input signals from poloidal array of pick-up coils, flux loops, and coil currents, and concatenates them into a $68$-element vector. EFITNN supports the prediction of $8$ scalar values including six plasma shape parameters as well as safe factor in 95% minor radius $q_{95}$ and plasma inductance $l_{i}$. Besides, it can produce predictions for poloidal magnetic flux distribution $\Psi_{\text{rz}}$ and toroidal current density distribution $J_{p}$ with a resolution up to $129\times 129$. For these heterogeneous outputs, EFITNN adopts a shared $4$-layer MLPs. Subsequently, for scalar prediction, another $4$-layer MLPs with residual connections [54] are leveraged, while $\Psi_{\text{rz}}$ and $J_{p}$ relies on deconvolutional networks [74]. Note that each MLP layer consists of $512$ neutrons with a GeLU activation function. EFITNN adopts simultaneous training for this multi-task learning, and we kindly refer to [38] for the training details. Fig. 8 manifests the prediction accuracy after training. In addition, if only plasma shape and position parameters are predicted, EFITNN, without the use of deconvolutional networks, consumes $0.08$ ms on NVIDIA A100 Tensor Core GPU with TensorRT, fully supporting the real-time deployment.
4.3.3 DNN Structure for PPO
We train the RL agent using PPO [53], which, as an actor-critic framework, updates both the policy and value networks through the interactions with the well-trained dynamics model. In our experiments, both networks are structured as three-layer MLPs. The policy network uses a tanh activation function and consists of $64$ neurons per layer, while the value network employs a ReLU activation function and comprises $128$ neurons per layer.
4.4 Training of High-Fidelity Dynamics Models
4.4.1 Scheduled Sampling
[…]
Fig. 12: Schematic illustration of scheduled sampling. a, Without any preventive training measures, the autoregressive nature of the dynamics models implies a gradually larger deviation between the autoregressive output (blue line) and the true data (orange line), due to the compounding error. b, The scheduling sampling technique exposes the model to real data beyond the model output, with a probabilistic teacher-forcing ratio $p$, during the training.
Without any preventive training measures, the autoregressive nature of dynamics models leads to a gradually increasing deviation between the actual data and model predictions due to compounding error, as illustrated in Fig. 12a. Directly applying such a model in practice can result in exposure bias [75], incurring reduced efficacy and severe misalignment between the training and testing modalities. Therefore, during our training, we incorporate scheduled sampling [76, 77] to stochastically determine the input, by selecting between the model’s past predictive outputs and the current-step actual data according to a teacher-forcing ratio $p\in[0,1]$. In other words, as the ratio $p$ approaches $1$, a greater proportion of the real input data will be used. Conversely, as $p$ turns to 00, the model transitions to a strictly autoregressive mode, solely relying on the preceding steps’ predictions. During the training, we initialize the ratio $p$ to $1$ and linearly decay it to 00 over $300$ epochs. This approach ensures rapid convergence during training and facilitates a seamless shift towards an autoregressive scheme for training RL.
[…]
Fig. 13: The advantage of scheduled sampling over a noise layer. SS: Scheduled Sampling; LSTM w/ $(0.1\sigma$/$0.2\sigma$/$0.3\sigma)$ represents the combination of LSTM and a zero-mean Gaussian noise layer with a standard deviation of $0.1\sigma$/$0.2\sigma$/$0.3\sigma$.
In contrast, [11] adopts a different approach to intentionally add a zero-mean Gaussian noise layer to the training data. Nevertheless, apart from the conflict with the data smoothing operation during training, the comparison in Fig. 13 also suggests its inferiority to scheduled sampling. Specifically, the diverse impact on high-dimensional multi-task learning makes selecting the appropriate variance of added Gaussian noise challenging. For example, with an amplitude of $0.1\sigma_{x}$, where $\sigma_{x}$ indicates the standard deviation of the input data, the model exhibits the lowest error in predicting $\kappa$ but the highest error in predicting $I_{p}$. Conversely, the incorporation of scheduled sampling maintains stability in significantly reducing prediction errors for $25$ variables. Therefore, we primarily take account of scheduled sampling in this paper.
4.4.2 Adaptive Weight Adjustment
To unanimously provide multi-task predictions for $25$ variables, in Table 2, the calibration of relative weights assigned to the predictive losses of individual variables could heavily affect the model performance. Mathematically, for a space $\mathcal{V}$ containing $25$ variables, the aggregated loss function given the $44$-dimensional input $\boldsymbol{x}$ shall be a scalar $\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)$ computed from $25$ individual loss functions $L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)$ weighted by $c_{v}$, where $y_{v}$ and $\hat{y}_{v}$ correspond to actual data and model prediction for the variable $v\in\mathcal{V}$, respectively while $\boldsymbol{y}_{\mathcal{V}}$ and $\hat{\boldsymbol{y}}_{\mathcal{V}}$ denote the concatenation of $y_{v}$ and $\hat{y}_{v}$, $\forall v\in\mathcal{V}$. Due to the discrepancies in distribution and noise among the variables, a straightforward averaging of individual loss functions can adversely impact the model’s accuracy. Meanwhile, manual weight tuning is an arduous and time-intensive effort. Therefore, we utilize an adaptive weight adjustment technique by regarding the loss weight $c_{v}$ as a trainable single variate[45]. To prevent the loss weights from taking negative values during the optimization process, which could lead to negative loss values, we also employ a non-negative regularization constraint [44] to enhance the learning robustness. To sum up, the final loss function for training can be calculated as
$\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)=\sum_{v\in\mathcal{V}}\left[\frac{1}{2\cdot c_{v}^{2}}\cdot L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)+\ln\left(1+c_{v}^{2}\right)\right].$
4.4.3 Training Details
In training the dynamics model, we initially segment all the discharge data of training shots into consecutive groups with a fixed length of $30$ ms, followed by a shuffling process on all the segmented data. We use DataLoader, a PyTorch library, to load the processed training data, with a batch size set as $4,096$. In each training round, we randomly utilize a batch of data without replacement for training and an epoch completes until all groups have been utilized. We employ the Adam optimizer for optimizing network parameters, with an initial learning rate of $0.001$ for both the dynamics model and the adaptive weight adjustment. To prevent training instability due to an excessively high learning rate in the later stages of training, we adjust the learning rate using an exponential decay method. Specifically, we reduce the learning rate by a decay factor of $0.95$ every $5$ epochs, over a total of $300$ training epochs.
Finally, in addition to Fig. 3, Fig. 14 shows the accuracy for predicting all the $18$ coil currents.
Fig. 14: Goodness-of-fit of the dynamics model for coil currents in terms of the coefficient of determination $R^{2}$.
4.5 Training and Deployment of RL-based Magnetic Controller
4.5.1 State, Action, and Reward of RL
In this study, we utilize RL to control the plasma current ($I_{p}$) and six shape and position parameters ($a$, $\kappa$, $\delta_{l}$, $\delta_{u}$, $R$, $Z$), and denote the concatenation of current observations as $\boldsymbol{o}$. To ensure the integrity and stability of HL-3 plasma discharges, we select the actual waveform of shot #6226, an H-mode discharge with ELMs and a plasma current of $300$ kA, as the desired target $\boldsymbol{d}$ for control. As shown in Table 3, we consider a $14$-dimensional state $\boldsymbol{s}$ by amalgamating the observations $\boldsymbol{o}$ and desired target $\boldsymbol{d}$, while the $17$-dimensional action $\boldsymbol{a}$ encompasses voltage commands of $1$ CS coil and $16$ PF coils. To ensure control safety and hardware integrity, we impose operational constraints by stipulating that the dynamic rate of change between two consecutive voltage commands must not exceed $20$ V/ms. In other words, for any coil, any voltage command that changes beyond this limit will be clipped to adhere to the safety constraints.
Table 3: State and action space for RL-based magnetic control.
Name Description
State Space $\mathcal{S}$ Target and currently observed plasma current, and six shape and position parameters
($14$ dimensions)
Action Space $\mathcal{A}$ Voltage commands of CS and $8$ pairs of PF coils
($17$ dimension)
On the other hand, we devise the reward function $r(\boldsymbol{s},\boldsymbol{a})$ from two orthogonal perspectives. First, consistent with [9, 10], the reward function primarily takes account of the element-wise divergence between the current observation $\boldsymbol{o}\triangleq[o^{(1)},\cdots,o^{(7)}]$ and desired target $\boldsymbol{d}\triangleq[d^{(1)},\cdots,d^{(7)}]$. To merge the divergence components into a scalar reward and emphasize those control variables with sub-optimal control results, a smooth-max function [10] is leveraged. Second, owing to the constraints of the power module, the coil current is permitted to cross zero only once. Therefore, contingent on the coil current predicted by the dynamics model, we impose a penalty term to inhibit the agent from exploring control commands that would result in multiple zero crossings of the coil current. To sum up, the reward for a state-action pair $(\boldsymbol{s},\boldsymbol{a})$ is given as
$r(\boldsymbol{s},\boldsymbol{a};\boldsymbol{\omega},\alpha)=-\frac{\sum_{i=1}^{7}\omega_{i}|o^{(i)}-d^{(i)}|\exp(-\alpha|o^{(i)}-d^{(i)}|)}{\sum_{i=1}^{7}\omega_{i}\exp(-\alpha|o^{(i)}-d^{(i)}|)}+\sum\nolimits_{i=1}^{17}p_{i},$ (1)
where weights $\boldsymbol{\omega}\triangleq[\omega_{1},\cdots,\omega_{7}]$ in the softmax function provide the (relative) importance of each component and are initialized as $[3,2,2,3,1,2,2]$. The modulation factor $\alpha$, which affects the trade-offs between “easy” and “hard” (to satisfy) components [10], is set as $-1$ in our experiment. Besides, for $17$ coils, the penalty term $p_{i}$ for coil $i$ is designed as $-1$ when the coil current exceeds one zero-crossing while nulls otherwise.
4.5.2 Dynamics Model-based Training
We utilize the trained dynamics model as the interaction environment, through which the RL agent explores suitable actions to achieve the desired plasma current and shape. Specifically, as the HL-3 tokamak is a preliminarily established device and initiates the very first-round global open campaign this year, a significant number of discharge shots have been scheduled for international joint physical experiments. Consequently, the proposal for RL-based control has to share discharges with other experiments. In addition, given the potential uncertainty in freshly deploying RL-based control on HL-3, the early phase of a discharge, prior to $1500$ ms, is primarily allocated to other physics experiments. Therefore, during training, RL assumes the $1$-kHz control since the $1500$ ms of the discharge, and the initial observations $\boldsymbol{o}$ are synchronized with those of shot #6226, ensuring a smooth transition between PID control and RL control. Furthermore, we take episodic training from $1500$ ms to $2000$ ms of the discharge. Without loss of generality, at each time step $t$, the RL agent determines an action $\boldsymbol{a}_{t}$ according to the state $\boldsymbol{s}_{t}$, while the dynamics model yields next-step plasma current as well as the six shape and position parameters $\boldsymbol{o}_{t+1}$ based on $\boldsymbol{s}_{t}$ and $\boldsymbol{a}_{t}$. Note that given the slow-changing nature of the toroidal magnetic field and loop voltage, these two values are considered invariant and maintain the values same as the handover time of shot #6226. Hence, a reward $r(\boldsymbol{s}_{t},\boldsymbol{a}_{t})$ can be computed as in Equation (1). Finally, the episodic reward is derived from the discounted accumulative reward with a discount factor $\gamma$ set as $0.98$. To strike a balance between the variance and bias in the value function estimation, we employ the Generalized Advantage Estimation (GAE) [78] to compute the advantage function. In addition, other parameters for the PPO are set in a standard way. Specifically, we set the entropy coefficient, the clipping parameter $\epsilon$, and the $\lambda$ parameter for GAE as $0.05$, $0.2$, and $0.95$, respectively. Following this, the PPO algorithm [53] can iteratively update the agent’s actor and critic networks. Fig. 7a illustrates the tracking accuracy of the trained agent toward the target waveform within the simulation setting and suggests that the RL agent, supported by a data-driven dynamics model, has effectively explored and assimilated strategies that are beneficial for achieving discharge control objectives.
4.5.3 On-Device Deployment
To ascertain the practicability of the trained RL agent as well as the accuracy of the dynamics model, we integrate the PPO actor into the PCS of HL-3 to undertake RL-based control in real discharges. Since the PCS operates under the Redhat $6.9$ system and performs real-time computations in C language, we convert the Pytorch-based PPO actor into a compatible ONNX (Open Neural Network Exchange) [79] format, which offers cross-platform deployment capability and high execution efficiency. During real discharges, the RL agent receives plasma shape and position parameters from EFITNN [38] while the plasma current is measured by the CODIS-RTC system via reflective memory in real time. The agent then generates coil voltage commands and delivers them to actuators through the PCS, completing a closed-loop control period of $1$ ms, synchronized with that for the PID control.
As mentioned earlier, the first global open campaign on HL-3 has attracted a considerable number of international joint experiments, limiting the opportunities for testifying RL-based control. Eventually, in a shared discharge, the handover to RL-based control has to be delayed until other physics experiments involving auxiliary heating are completed. To avoid any potential plasma disturbances caused by shutting down the auxiliary heating at $1500$ ms, which could impact the analysis of experimental results, we have further postponed the start of RL-based control to $1600$ ms. We initially begin with a duration of $15$ ms for RL-based control and then revert to PID control. Afterward, we gradually extend the duration upon each successful execution, and ultimately, after $6$ effective shots, the experiment ends with a control duration of $100$ ms, with the last remaining shot attempting to implement a zero-shot control towards changed triangularity targets. Notably, to ensure stability incurred during these mode switches, we establish a $10$-ms transition period, during which a weighted sum of commands from both modes is undertaken and the weight of superseding control mode linearly increases. The limited opportunities hinder the on-device experiment for a longer RL-based takeover in practice.
4.5.4 Post-Experiment Analyses
As illustrated in Fig. 7b and Fig. 9a, the post-experiment replay for shot #$6691$ and shot #$6698$ is calculated using the dynamics model. Specifically, the dynamics model starts to evolve according to the input variables in Table 2, set as practical values at the handover time, as well as the practically executed coil voltage commands, which are generated by RL during on-device deployment. In other words, such a post-experiment replay represents the autoregressive output of the dynamics model, and its closeness to the EFITNN-based diagnostic measurement contributes to verifying the accuracy of the dynamic model.
On the other hand, as shown in Fig. 4d and Fig. 9b, to visually display the prediction and control effect on the plasma shape, we calculate the four extreme points of the cross-section based on the six shape parameters obtained from the EFITNN computation. The plasma boundary is then approximated as the connection of four-quarter ellipses [80].
Acknowledgments
The authors would like to thank the entire HL-3 team for providing experimental data and engineering support for the verification experiments in the HL-3 tokamak. This work was in part supported by the National Nature Science Foundation of China under Grant No. U21A20440.
Author contributions
N.N.W., Z.Y., R.L., and N.W. contributed equally to this work. R.L., M.X., Z.Z., and W.Z. conceived this collaboration. N.N.W., R.L., and N.W. led the design & implementation of the dynamics models and reinforcement learning approach; Z.Y. and G.Z. designed and implemented EFITNN; Y.C., X.G., F.G., and B.L. contributed to the real-time on-device implementation. N.N.W, Z.Y., R.L., and N.W. discussed the results; Q.D. and J.L. contributed to the analyses and illustration of some results. N.N.W. and R.L. wrote this manuscript; Z.Y. and N.W. revised the manuscript; all other authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.
Data availability
Data sets generated during the current study are available from the corresponding author on request.
Code availability
Codes used during the current study are available from the corresponding author on request.